Draw a triangle with vertices (-6,3), (4,13), and (10,-5). It tells me to find the slope intercept and general form equations for the three sides of the triangle. So I found the slope-intercept form equations for three sides, which are Y=X + 9, Y=-1/2X+0 and y=-3x+25. Does the general form equations just mean standard form? AKA x-y=-9 instead?
Answer by rothauserc(4718)

the general form for the equation of a line is
:
Ax +By +C = 0, where A, B, C are real numbers, also A or B can be zero, but not both at the same time.
:
therefore, your three lines have the general form
:
x -y +9 = 0
:
x +2y = 0
:
3x +y -25 = 0
: